solve the following system of equations using substitution.
-2x-10y=-12
3x+y=4

Answer by mananth(16946)

3 x + 1 y = 4 .............1
y = 4-3x

-2 x + -10 y =-12 .....................2
Substitute y in (2)
-2x-10(4-3x)= -12
-2x-40 +30x= -12
-2x+30 x = 28
28 x = 28
/ 28
x= 1
Plug the value of x in (1)
3 x + 1 y = 4
3 * 1 + 1 y = 4
3 + 1 y = 4
1 y = 1
/ 1
y= 1
Solution (1,1)
m.ananth@hotmail.ca